Calculate the pH of: A) 12M hydrochloric acid, HCl. B) 0.40 M nitric acid, HNO3.
What is the pH of 0.10 M hypochlorous acid (HOCl)
Answer #1

HCl is strong acid

HCl(aq) --------------> H^+ (aq) + Cl^- (aq)'

12M                            12M

[H^+]     =   [HCl]

[H^+]   = 12M

PH    = -log[H^+]

          = -log12

         = -1.08

b.    HNO3(aq) -------------> H^+ (aq) + NO3^- (aq)

      0.40M                          0.40M

   [H^+]    = [HNO3]

[H^+]   = 0.40M

PH   = -log[H^+]

          = -log0.40

            = 0.3979

HOCl is weak acid

        HOCl(aq) ---------------> H^+ (aq) + OCl^- (aq)

I      0.1                                 0               0

C        -x                                 +x              +x

E       0.1-x                                +x              +x

              Ka    = [H^+][OCl^-]/[HOCl]

             2.8*10^-8   = x*x/(0.1-x)

            2.8*10^-8*(0.1-x) = x^2

             x   = 5.3*10^-5

          [H^+]   = x   = 5.3*10^-5M

         PH   = -log[H^+]

                    = -log(5.3*10^-5)

                     = 4.2757 >>>>answer